OpenStudy (anonymous):

Will Medal And Medal. Solve: 52 - 2x = 4 A. -28 B. -24 C. 24 D. 28

OpenStudy (parker.goodbar):

do you need work with it

OpenStudy (parker.goodbar):

its C

OpenStudy (parker.goodbar):

the reason is that you need to move the -2x over so add 2x to both sides then subtract 4 into the 52 to get 48 then divide the 2 into the 48 and get 24

OpenStudy (kakarot):

x = 24

OpenStudy (parker.goodbar):

or 52-2x=4 +2x +2x 52=4+2x -4 -4 48=2x /2 /2 24=x

OpenStudy (anonymous):

so always do the opposit of the sign

OpenStudy (parker.goodbar):

yes

OpenStudy (parker.goodbar):

so it will cancel the # an example is if i have a 1+5x=6 i would subtract 5x to both sides

OpenStudy (parker.goodbar):

does that answer your question